2. An investment produces a cash flow of $420 in year 1, $137 in year 2, and $797 in year 3. If the required rate of return is 15%, what price should you pay for the investment at time zero?
Answer #1

Answer: $ 992.85

Year Cash Flows PV of Cash Flows Present Value
1 $ 420 ( 1/1.15) = 0.8696 $ 365.23
2 137 (1 / 1.15) 2 = 0.7561 103.59
3 797 ( 1 / 1.15 ) 3 = 0.6575 524.03
$ 992.85